OpenStudy (anonymous):

Find an equation for the line satisfying the given conditions. Through (5, 4) and parallel to 4x - 3y = 7.

OpenStudy (anonymous):

Whenever you see parallel right away you know it has the same slope

OpenStudy (anonymous):

first solve for this by rewriting the equation in slope intercept form

OpenStudy (anonymous):

once you have the slope, plug in the point to y=mx+b and solve for b. This will give you the equation!!

Directrix (directrix):

4x - 3y = 7 Do the algebra to solve for y: y = (4/3)x - (7/3) --> y = mx +b form of a linear equation where m is the slope and b is the y-intercept. The slope of the line is (4/3). The question now becomes the following: What is the equation of the line with slope (4/3) and passing through the point (5,4) Use. the point-slope form of a linear equation, (y - y1) = m(x - x1) where (x1,y1) is a point on the line and m is the slope of the line. y -4 = (4/3) (x - 5) Solving for y to get the equation in slope-intercept form, y = (4/3) x - (8/3) --> Check the work. Thanks. @orbie

OpenStudy (anonymous):

It's right but how did you get -(8/3)? @Directrix

Directrix (directrix):

@orbie --> Did you understand how I got this: y -4 = (4/3) (x - 5) ?

OpenStudy (anonymous):

@Directrix yes

Directrix (directrix):

y -4 = (4/3) (x - 5) y - 4 = (4/3)x + (4/3)(-5) y - 4 = (4/3)x - 20/3 y = (4/3)x - 20/3 + 4 y = (4/3)x - 20/3 + 12/3 y = (4/3)x - (8/3)

OpenStudy (anonymous):

ooh okay thank you!